# Counterfactual Uncertainty Quantification of Factual Estimand of Efficacy from Before-and-After Treatment Repeated Measures Randomized Controlled Trials

### v3

@@ 4.5 Estimated average $Rx:C$ efficacy is unbiased
Population effect of treatment $Rx$ or $C$ is defined as response $Y$ averaged over infinitely many patients each with their value $Z$. In the population space, what is of interest is the value $\mathbbm{E}[Y]$ at $\mathbbm{E}[Z]$.
@@
so there is bias if $c>0$.
@@ 5 Guidelines for applying the ETZ modeling principle
ETZ is a modeling principle that enables counterfactual uncertainty quantification (CUQ) for randomized studies with Before-and-After Repeated Measures (BAtRM). Scientists faced with highly variable outcome measure (i.e., large measurement error) situations can benefit from this principle, so long as they can extract the three inputs needed by ETZ, working out details for proper CUQ in their settings themselves.
For example, response to therapies for amyotrophic lateral sclerosis (ALS) is highly variable. Outcome measured on the ALS Functional Rating Scale–Revised (ALSFRS-R) scale (FDA (2021)) can be modeled with a random coefficient BAtRM model (e.g., using the RANDOM statement in Proc Mixed of SAS), providing inputs needed by ETZ.
Scientists can apply ETZ to efficacy measured as a ratio, which is common for binary and time-to-event outcome measures, provided they are aware of the following:
@@
RCTs in therapeutic areas such as immunology often are BAtRM studies with binary endpoints. For example, Morand et al. (2023) reported result of the Phase 3 study SLE-BRAVE-I, testing baricitinib for the treatment of systemic lupus erythematosus (SLE). SLE-BRAVE-I was a double-blind, randomized, placebo-controlled, parallel-group Phase 3 study. A total of 760 participants were randomly assigned to receive at least one dose of baricitinib 4 mg (n=252), baricitinib 2 mg (n=255), or placebo (n=253). Primary endpoint was the proportion of patients achieving an SLE Responder Index (SRI)-4 response at week 52 in the baricitinib 4 mg treatment group compared to the placebo group. Achieving SRI-4 is a binary outcome, a composite responder index that includes improvement in disease activity (defined as at least a 4-point improvement in the SLEDAI-2K score) without worsening of the overall condition (no increase of $\geq$ 0.3 points [10 mm] in the Physician Global Assessment from baseline) or the development of significant disease activity in new organ systems (no new BILAG A or more than one new BILAG B).
To develop treatments for diseases such as ALS (which affects an estimated 31,000 patients in the U.S.) and SLE (which affects approximately 200,000 patients in the U.S.), variability in the outcome measure is a known challenge. The CUQ principle for reducing uncertainty quantification is developed to help meet this challenge, for the greater good.
@@ 6 Key messages
ETZ is an uncertainty quantification (UQ) principle for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures (BAtRM). Conducting RCT with BAtRM is standard practice in therapeutic areas affecting more than half a billion patients worldwide. By modeling individual patient outcomes in BAtRM studies, ETZ enables counterfactual uncertainty quantification (CUQ). CUQ is typically smaller than current factual uncertainty quantification. Reduction in uncertainty quantification translates to reduction in sample size. Statistical or digital counterfactual point estimation involving a biomarker with measurement error is unbiased for efficacy in the entire population but biased in subgroups.
7 Acknowledgement

### v4

@@ 5.5 Estimated average $Rx{:}C$ efficacy is unbiased
Population effect of treatment $Rx$ or $C$ is defined as response $Y$ averaged over infinitely many patients, each with their value $Z$. In the population space, what is of interest is the value $\mathbbm{E}[Y]$ at $\mathbbm{E}[Z]$.
@@
so there is bias if $c>0$. That would happen if the target patient population in the actual RCT does not align with the patient population in the learning study. Table 5 advises extraordinary caution in this situation using two examples.
MMSE scoring
Study Mild/Early Moderate
EXPEDITION1 $\checkmark$ $\checkmark$
Start of EXPEDITION2 $\checkmark$ $\checkmark$
End of EXPEDITION2 $\checkmark$
EXPEDITION3 $\checkmark$
Training ProCovA $\checkmark$
Applying ProCovA $\checkmark$ $\checkmark$
Table 5: From Section 5.4, and experience from EXPEDITION1/2/3 (which did not meet their primary objectives), extraordinary caution is advisable in predicting potential outcome when the predictor contains measurement error and the learning study patient population differs from the application study patient population. EXPEDITIONS 1, 2, and 3 staged Alzheimer’s disease patients by the Mini–Mental State Examination (MMSE) score, with an MMSE score of 20 to 26 classified as mild, while a score of 16 to 19 is moderate. From Honig et al. (2018), MMSE has an estimated $Var(E)$ of 2.06 which is not ignorable since the estimated MMSE difference between solanezumab and the placebo is 0.49. EXPEDITION 2 started by targeting mild and moderate patients, but changed to targeting mild patients after learning from EXPEDITION1’s result. EXPEDITION3 targeted only mild patients after learning from data on mild patients pooled from EXPEDITION1 and EXPEDITION2. From Schuler et al. (2022), it appears ProCovA was applied to mild and moderate patients with MMSE between 14 and 26 in the DHA supplement study, while the algorithm appears to have been trained on early stage patients.
@@ 6 Guidelines for applying the ETZ modeling principle
ETZ is a modeling principle that enables Counterfactual Uncertainty Quantification (CUQ) for randomized studies with Before-and-After Repeated Measures (BAtRM). Scientists faced with highly variable outcome measure (i.e., large measurement error) situations can benefit from this principle, provided that they can extract the three inputs needed by ETZ, working out details for proper CUQ in their settings themselves.
For example, highly variable responses to therapies for amyotrophic lateral sclerosis (ALS, affecting $\approx$ 31,000 patients in the U.S.), measured on the ALS Functional Rating Scale–Revised (ALSFRS-R) scale (FDA (2021)) can be modeled with a random coefficient BAtRM model (e.g., using the RANDOM statement in Proc Mixed of SAS), providing inputs needed by ETZ. As another example, RCTs in immunology such as for systemic lupus erythematosus (SLE, affecting $\approx$ 200,000 patients in the U.S.), often are BAtRM studies with binary endpoints. For binary and time-to-event outcomes, efficacy is often measured as a ratio. The CUQ principle for reducing uncertainty quantification can be applied to ratio efficacy, provided one is aware of the following:
@@ 7 Key messages
ETZ is an uncertainty quantification (UQ) principle for Randomized Controlled Trials (RCTs) with Before-and-After treatment Repeated Measures (BAtRM). Conducting RCT with BAtRM is standard practice in therapeutic areas affecting more than half a billion patients worldwide. By modeling individual patient outcomes in BAtRM studies, ETZ enables counterfactual uncertainty quantification (CUQ). CUQ is typically smaller than current factual uncertainty quantification. Statistical or digital counterfactual point estimation involving a biomarker with measurement error is unbiased for efficacy in the entire population but biased in subgroups.
8 Acknowledgments
@@
We thank the AE for on-point comments, particularly for requesting us to comment on ProCovA.
@@ Appendix A Glossary of terminologies
Appendix B Details of CUQ variance reduction calculations
Detailed calculations of the two-step CUQ process that led to the pie chart for EXPEDITION3 in Figure 3 are given here. Calculations for the AbbVie data are similar.
Table 4 presents summary statistics for EXPEDITION3 along with their ETZ variance components. Detailed calculation showing baselining reduces variance in milestone visit by 43.02 which is 31.77% of total variance is as follows.
$Var(Y^{[m]})$ $=$ $Var(Z)+Var(Trt)+Var(E)$
$=$ 135.39135.39 135.39135.39
$Var(Y^{[change]})$ $=$ $Var(Trt)+2\times Var(E)$
$=$ 92.3792.37 92.3792.37
$Var(Y^{[m]})-Var(Y^{[change]})$ $=$ $Var(Z)-Var(E)$
$=$ $53.802-10.778$
$=$ 43.0243.02 43.0243.02
$\frac{Var(Y^{[m]})-Var(Y^{[change]})}{Var(Y^{[m]})}$ $=$ $31.77\%.$
Detailed calculation showing the counterfactual aspect of CUQ further reduces variance by 21.56 beyond baselining by 11.67% is as follows.
$Var(Y^{f})$ $=$ $2\times Var(Trt)+4\times Var(E)$
$=$ $2\times 70.809+4\times 10.778$
$=$ 184.73184.73 184.73184.73
$Var(Y^{cf})$ $=$ $2\times Var(Trt)+2\times Var(E)$
$=$ $2\times 70.809+2\times 10.778$
$=$ 163.174163.174 163.174163.174
$Var(Y^{f})-Var(Y^{cf})$ $=$ $2\times Var(E)$
$=$ $184.73-163.174$
$=$ 21.55621.556 21.55621.556
$\frac{Var(Y^{f})-Var(Y^{cf})}{Var(Y^{f})}$ $=$ $11.67\%.$
Hence, CUQ reduces
$31.77\%+11.67\%\times(1-31.77\%)=39.73\%$ (34)
of the total variance in EXPEDITION3.